### v1

@@ Abstract
One-dimensional nuclear magnetic resonance (NMR) spectroscopy is essential for molecular structure elucidation in organic synthesis, drug discovery, natural product characterization, and metabolomics, yet its interpretation remains heavily dependent on expert knowledge and difficult to scale. Although machine learning has been applied to NMR spectrum prediction, library retrieval, and structure generation, these tasks have evolved in isolation using simulated data and incompatible spectral representations, limiting their utility under real experimental scenarios. Here we present NMRPeak, a unified cross-modal learning system that integrates these three tasks through experimentally grounded design. We curate approximately 1.8 million experimental and simulated spectra to construct the largest benchmark for NMR-based structure elucidation and systematically quantify the distribution shift between these domains. We introduce a chemically-aware adaptive tokenizer that dynamically balances discretization granularity to preserve spectral semantics while controlling vocabulary size, and an assignment-free peak-aware similarity metric that enables direct comparison between predicted and experimental spectra. Through a unified molecule-to-spectrum paradigm and synergistic coupling of prediction, retrieval, and generation modules, NMRPeak achieves transformative performance on experimental benchmarks: it overcomes the longstanding simulation-to-experiment gap in spectrum prediction while delivering over 95% top-1 accuracy in molecular retrieval and approximately 75% top-1 accuracy in stereochemistry-aware de novo structure generation. These capabilities enable rapid validation of synthetic products, characterization of unknown natural products, metabolite identification in complex biological samples, and accelerated lead compound optimization, establishing a foundation for automated, high-throughput molecular structure elucidation in organic synthesis, drug discovery, and chemical biology.
@@ 1 Introduction
One-dimensional nuclear magnetic resonance (NMR) spectroscopy provides fingerprint-level information that reflects the local chemical environments and connectivity of atoms in organic molecules [hu2023machine]. As a result, it remains one of the most widely used experimental techniques for molecular structure elucidation. However, conventional NMR interpretation relies heavily on expert knowledge and manual reasoning, making the process time-consuming, labor-intensive, and difficult to scale to complex structures or high-throughput experimental pipelines [xue2023advances, kuhn2024nuclear, das2025exploring]. Artificial intelligence (AI) methods have been increasingly applied to address these challenges in NMR-based analysis. Conceptually, NMR interpretation can be formulated as a pair of complementary problems: the forward problem predicts NMR spectra from molecular structures, while the inverse problem infers molecular structures from observed spectra. In practice, AI-assisted NMR analysis has evolved into three major task paradigms: Prediction models learn to simulate spectra from molecular structures, enabling rapid computational screening and spectrum validation; Retrieval methods search for candidate molecules in databases given a query spectrum, leveraging pre-existing chemical knowledge for efficient structure identification; and Generation approaches attempt to directly construct molecular structures from spectral inputs, enabling de novo structure elucidation when database matches are unavailable.
@@
First, spectral representations are misaligned across tasks. Prediction models typically rely on atom-level assignments that map individual atoms to chemical shifts, requiring fully annotated training data that are rarely available in experimental settings. In contrast, retrieval and generation methods often operate on unassigned experimental spectra represented as global peak sets, where equivalent atoms are merged and additional information such as splitting patterns, coupling constants and peak integrals in 1H NMR is present. This representational divide creates a critical gap: atom–shift representations, while computationally convenient, cannot be directly applied to real experimental spectra, whereas molecule–spectrum representations naturally align with how spectra are acquired and reported in practice. Compounding this challenge, spectral discretization strategies remain inconsistent and fundamentally unresolved. Continuous spectral data points preserve raw spectral information but they are highly sensitive to preprocessing choices such as interpolation, resolution, phase correction, and truncation. Discrete peak list representations align with experimental reporting but introduce a critical trade-off in discretization granularity during model tokenization. Fine-grained binning (e.g., 0.1 ppm resolution for 13C NMR shifts yielding 2,000 bins in Wang et al. [yao2023conditional, xue2025nmrmind]) leads to large vocabularies and severe data sparsity. Conversely, coarse-grained binning (e.g., fewer than 100 bins in Hu et al. [hu2024accurate, hu2025pushing]) collapses distinct spectra into nearly identical representations, obscuring chemically meaningful differences. This divergence highlights the absence of a principled, chemically informed solution to the representation problem.
@@ 2.3 NMRPeak-R: Cross-modal retrieval between molecules and NMR spectra
Figure 6: Contribution of individual similarity metrics of NMRPeak-R. a–f, Impact of weight variations for each component on retrieval accuracy across diverse configurations, with data points and error bars representing the mean and variance, respectively. For all panels, SME, SSE, and SSR denote spectrum-to-molecule embedding, spectrum-to-spectrum embedding, and spectrum-to-spectrum rule-based similarity, respectively. C and H correspond to 13C and 1H NMR peaks, and F denotes molecular formula constraints.
To dissect the contribution of each component, Fig. 5a compares multiple retrieval variants. Introducing NMRPeak-R-S2M-SSE yields measurable improvements over NMRPeak-R-S2M-SME, confirming the utility of predicted spectra as intermediate representations. Incorporating NMRPeak-R-S2M-SSR leads to a more pronounced performance gain, closely approaching that of the full combined model. These results highlight the importance of directly modeling spectral physics rather than relying exclusively on latent semantic alignment. Further analysis of weighting strategies (see Fig. 5b and Fig. 6) reveals a consistent pattern: overemphasizing embedding-based similarity degrades retrieval performance, whereas increasing the contribution of explicit peak-aware rules steadily improves accuracy. This observation suggests a general retrieval principle in scientific domains that latent embeddings are effective for candidate recall while explicit domain rules are essential for fine-grained discrimination.
@@ 2.4 NMRPeak-G: End-to-end molecular structure generation from NMR spectra
Figure 8: Ablation study and weighting analysis of NMRPeak-G. a, Comparison of structural inference accuracies for different generative variants across various input modalities. The bar charts evaluate the performance of the MST baseline, NMRPeak-G-Raw, and various reranking configurations (SME, SSE, SSR, and their combinations) for top-1, top-5, and top-10 candidates. b, Ternary plots illustrating the impact of weighting strategies on Ttop-1 generation accuracy for diverse configurations. For all panels, SME, SSE, and SSR denote spectrum-to-molecule embedding, spectrum-to-spectrum embedding, and spectrum-to-spectrum rule-based similarity, respectively. C and H correspond to 13C and 1H NMR peaks, and F denotes molecular formula constraints.
Figure 9: Contribution of individual similarity metrics of NMRPeak-G. a–f, Impact of weight variations for each component on generation accuracy across diverse configurations, with data points and error bars representing the mean and variance, respectively. For all panels, SME, SSE, and SSR denote spectrum-to-molecule embedding, spectrum-to-spectrum embedding, and spectrum-to-spectrum rule-based similarity, respectively. C and H correspond to 13C and 1H NMR peaks, and F denotes molecular formula constraints.
@@
Ablation studies presented in Fig. 8 and Fig. 9 further elucidate the contributions of individual components. Specifically, NMRPeak-G-Rerank-Base, by incorporating chemical validation, yields a tangible improvement over NMRPeak-G-Raw baseline, while NMRPeak-G-Rerank-Combine leverages multi-dimensional ranking strategies to further optimize the ranking of candidate molecules. These trends across different ranking components closely mirror the observations from NMRPeak-R, reinforcing a fundamental principle that emerges across both retrieval and generation tasks.
@@ 5.3.1 Problem formulation
Given two NMR spectra of the same type (either 1H NMR or 13C NMR), we aim to compute a normalized similarity score in the range $[0,1]$ that quantifies their structural resemblance. Each spectrum is represented as an unordered set of peaks without atom-level assignments. For 1H NMR, each peak is characterized by a five-tuple: chemical shift range $\delta_{\min},\delta_{\max}]$, multiplicity, integration (number of hydrogens $n_{\text{H}}$), and coupling constants. For 13C NMR, each peak is represented solely by its chemical shift $\delta$. The similarity score enables direct evaluation of NMRPeak-P predictions and facilitates spectrum-to-molecule reranking in both NMRPeak-R and NMRPeak-G pipelines.
@@ 5.3.2 Two-stage peak matching procedure
Stage I: Global optimal matching. Without loss of generality, assume $|\mathcal{A}^{\prime}|\leq|\mathcal{B}^{\prime}|$. We construct a cost matrix $\mathbf{C}\in\mathbb{R}^{|\mathcal{A}^{\prime}|\times|\mathcal{B}^{\prime}|}$ for the linear assignment problem, with different formulations for 13C and 1H NMR.
@@ 5.3.3 Similarity scoring with multi-feature integration
$\ +0.05\cdot\mathbb{1}[\text{mult}(a_{\mu(i)})=\text{mult}(b_{\nu(j)})]$
$\ +0.1\cdot\mathbb{1}[n_{\text{H}}(a_{\mu(i)})=n_{\text{H}}(b_{\nu(j)})]$
@@
where $\mathbb{1}[\cdot]$ is the indicator function, $\text{mult}(\cdot)$ denotes peak multiplicity, and $s_{\text{coupling}}$ computes the similarity between sorted coupling constant lists:
@@ 5.4.1 Objective functions
Spectrum prediction loss (NMRPeak-P). NMRPeak-P is trained to autoregressively predict spectral tokens from molecular structure. The model takes as input molecular features encoded using the same protocol as Uni-Mol [zhouuni] (atomic elements and 3D coordinates when available), and generates a sequence of spectral tokens obtained from the chemistry-aware tokenizer. The model is trained with a standard autoregressive language modeling objective:
@@
Contrastive retrieval loss (NMRPeak-R). NMRPeak-R learns aligned representations of molecules and spectra through contrastive learning. The model encodes molecular features $\mathbf{m}$ into embedding $\mathbf{e}_{\text{mol}}$ and spectral tokens $\mathbf{z}$ (from the chemistry-aware tokenizer) into embedding $\mathbf{e}_{\text{spec}}$. For a batch of $B$ molecule-spectrum pairs, the model computes a similarity matrix scaled by temperature $\tau_{\text{cl}}$:
@@
where $\mathbf{E}_{\text{mol}}\in\mathbb{R}^{B\times d}$ and $\mathbf{E}_{\text{spec}}\in\mathbb{R}^{B\times d}$ are the normalized embedding matrices. The contrastive loss is computed bidirectionally:
@@
Molecular generation loss (NMRPeak-G). NMRPeak-G performs conditional molecule generation from spectral information. The model takes as input spectral tokens $\mathbf{z}$ and optionally molecular formula tokens $\mathbf{f}$ from the chemistry-aware adaptive tokenizer, and autoregressively generates SMILES tokens $\{s_{1},s_{2},\ldots,s_{L}\}$ obtained from Smirk-based tokenization [wadell2024smirk]:

### v2

@@ Abstract
One-dimensional nuclear magnetic resonance (NMR) spectroscopy is essential for molecular structure elucidation in organic synthesis, drug discovery, natural product characterization, and metabolomics, yet its interpretation remains heavily dependent on expert knowledge and difficult to scale. Although machine learning has been applied to NMR spectrum prediction, library retrieval, and structure generation, these tasks have evolved in isolation using simulated data and incompatible spectral representations, limiting their utility under real experimental scenarios. Here we present NMRPeak, a unified cross-modal learning system that integrates these three tasks through experimentally grounded design. We curate approximately 1.8 million experimental and simulated spectra to construct the largest benchmark for NMR-based structure elucidation and systematically quantify the distribution shift between these domains. We introduce a chemically-aware adaptive tokenizer that dynamically balances discretization granularity to preserve spectral semantics while controlling vocabulary size, and an assignment-free peak-aware similarity metric that enables direct comparison between predicted and experimental spectra. Through a unified molecule-to-spectrum paradigm and synergistic coupling of prediction, retrieval, and generation modules, NMRPeak achieves transformative performance on experimental benchmarks: it overcomes the longstanding simulation-to-experiment gap in spectrum prediction while delivering over 95% top-1 accuracy in molecular retrieval and approximately 75% top-1 accuracy in stereochemistry-aware de novo structure generation. These capabilities establish a foundation for automated, high-throughput molecular structure elucidation in organic synthesis, drug discovery, and chemical biology.
@@ 1 Introduction
One-dimensional nuclear magnetic resonance (NMR) spectroscopy provides fingerprint-level information that reflects the local chemical environments and connectivity of atoms in organic molecules [hu2023machine]. As a result, it remains one of the most widely used experimental techniques for molecular structure elucidation [hartman2016benchmark, lin2021unravelling, lin2022combining, lin2022machine, atwi2022automated, you2025decoding, fan2026relative]. However, conventional NMR interpretation relies heavily on expert knowledge and manual reasoning, making the process time-consuming, labor-intensive, and difficult to scale to complex structures or high-throughput experimental pipelines [xue2023advances, kuhn2024nuclear, das2025exploring, luo2025deep]. Artificial intelligence (AI) methods have been increasingly applied to address these challenges in NMR-based analysis. Conceptually, NMR interpretation can be formulated as a pair of complementary problems: the forward problem predicts NMR spectra from molecular structures, while the inverse problem infers molecular structures from observed spectra. In practice, AI-assisted NMR analysis has evolved into three major task paradigms: Prediction models learn to simulate spectra from molecular structures, enabling rapid computational screening and spectrum validation; Retrieval methods search for candidate molecules in databases given a query spectrum, leveraging pre-existing chemical knowledge for efficient structure identification; and Generation approaches attempt to directly construct molecular structures from spectral inputs, enabling de novo structure elucidation when database matches are unavailable.
@@
First, spectral representations are misaligned across tasks. Prediction models typically rely on atom-level assignments that map individual atoms to chemical shifts [paruzzo2018chemical, jonas2019rapid, kwon2020neural, guan2021real, zou2023deep, ai2024very, yan2025general, bhadauria2025cascade, xu2025toward], requiring fully annotated training data that are rarely available in experimental settings. In contrast, retrieval and generation methods often operate on unassigned experimental spectra represented as global peak sets, where equivalent atoms are merged and additional information such as splitting patterns, coupling constants and peak integrals in 1H NMR is present. This representational divide creates a critical gap: atom–shift representations, while computationally convenient, cannot be directly applied to real experimental spectra, whereas molecule–spectrum representations naturally align with how spectra are acquired and reported in practice. Compounding this challenge, spectral discretization strategies remain inconsistent and fundamentally unresolved. Continuous spectral data points preserve raw spectral information but they are highly sensitive to preprocessing choices such as interpolation, resolution, phase correction, and truncation. Discrete peak list representations align with experimental reporting but introduce a critical trade-off in discretization granularity during model tokenization. Fine-grained binning (e.g., 0.1 ppm resolution for 13C NMR shifts yielding 2,000 bins in Wang et al. [yao2023conditional, xue2025nmrmind]) leads to large vocabularies and severe data sparsity. Conversely, coarse-grained binning (e.g., fewer than 100 bins in Hu et al. [hu2024accurate, hu2025pushing]) collapses distinct spectra into nearly identical representations, obscuring chemically meaningful differences. This divergence highlights the absence of a principled, chemically informed solution to the representation problem.
@@ 2.3 NMRPeak-R: Cross-modal retrieval between molecules and NMR spectra
To dissect the contribution of each component, Fig. 5a compares multiple retrieval variants. Introducing NMRPeak-R-S2M-SSE yields measurable improvements over NMRPeak-R-S2M-SME, confirming the utility of predicted spectra as intermediate representations. Incorporating NMRPeak-R-S2M-SSR leads to a more pronounced performance gain, closely approaching that of the full combined model. These results highlight the importance of directly modeling spectral physics rather than relying exclusively on latent semantic alignment. Further analysis of weighting strategies (see Fig. 5b and Supplementary Figure 1) reveals a consistent pattern: overemphasizing embedding-based similarity degrades retrieval performance, whereas increasing the contribution of explicit peak-aware rules steadily improves accuracy. This observation suggests a general retrieval principle in scientific domains that latent embeddings are effective for candidate recall while explicit domain rules are essential for fine-grained discrimination.
@@ 2.4 NMRPeak-G: End-to-end molecular structure generation from NMR spectra
Figure 7: Ablation study and weighting analysis of NMRPeak-G. a, Comparison of structural inference accuracies for different generative variants across various input modalities. The bar charts evaluate the performance of the MST baseline, NMRPeak-G-Raw, and various re-ranking configurations (Base, SME, SSE, SSR, and their combinations) for top-1, top-5, and top-10 candidates. b, Ternary plots illustrating the impact of weighting strategies on top-1 generation accuracy for diverse configurations. For all panels, SME, SSE, and SSR denote spectrum-to-molecule embedding, spectrum-to-spectrum embedding, and spectrum-to-spectrum rule-based similarity, respectively. C and H correspond to 13C and 1H NMR peaks, and F denotes molecular formula constraints.
@@
Ablation studies presented in Fig. 7 and Supplementary Figure 2 further elucidate the contributions of individual components. Specifically, NMRPeak-G-Rerank-Base, by incorporating chemical validation, yields a tangible improvement over NMRPeak-G-Raw baseline, while NMRPeak-G-Rerank-Combine leverages multi-dimensional ranking strategies to further optimize the ranking of candidate molecules. These trends across different ranking components closely mirror the observations from NMRPeak-R, reinforcing a fundamental principle that emerges across both retrieval and generation tasks.
@@ 5.3.1 Problem formulation
Given two NMR spectra of the same type (either 1H NMR or 13C NMR), we aim to compute a normalized similarity score in the range $[0,1]$ that quantifies their structural resemblance. Each spectrum is represented as an unordered set of peaks without atom-level assignments. For 1H NMR, each peak is characterized by a five-tuple: chemical shift range $\delta_{\min},\delta_{\max}]$, multiplicity, integration (number of hydrogens $n_{\text{H}}$), and coupling constants. For 13C NMR, each peak is represented solely by its chemical shift $\delta$. The similarity score enables direct evaluation of NMRPeak-P predictions and facilitates spectrum-to-molecule re-ranking in both NMRPeak-R and NMRPeak-G pipelines.
@@ 5.3.2 Two-stage peak matching procedure
Stage I: Global optimal matching. Without loss of generality, assume $|\mathcal{A}^{\prime}|\leq|\mathcal{B}^{\prime}|$. We construct a cost matrix $\mathbf{C}\in\mathds{R}^{|\mathcal{A}^{\prime}|\times|\mathcal{B}^{\prime}|}$ for the linear assignment problem, with different formulations for 13C and 1H NMR.
@@ 5.3.3 Similarity scoring with multi-feature integration
$\ +0.05\cdot\mathds{1}[\text{mult}(a_{\mu(i)})=\text{mult}(b_{\nu(j)})]$
$\ +0.1\cdot\mathds{1}[n_{\text{H}}(a_{\mu(i)})=n_{\text{H}}(b_{\nu(j)})]$
@@
where $\mathds{1}[\cdot]$ is the indicator function, $\text{mult}(\cdot)$ denotes peak multiplicity, and $s_{\text{coupling}}$ computes the similarity between sorted coupling constant lists:
@@ 5.4.1 Learning objectives
Spectrum prediction loss of NMRPeak-P. NMRPeak-P is trained to autoregressively predict spectral tokens from molecular structure. The model takes as input molecular features encoded using the same protocol as Uni-Mol [zhouuni] (atomic elements and 3D coordinates when available), and generates a sequence of spectral tokens obtained from the chemistry-aware tokenizer. The model is trained with a standard autoregressive language modeling objective:
@@
Contrastive retrieval loss of NMRPeak-R. NMRPeak-R learns aligned representations of molecules and spectra through contrastive learning. The model encodes molecular features $\mathbf{m}$ into embedding $\mathbf{e}_{\text{mol}}$ and spectral tokens $\mathbf{z}$ (from the chemistry-aware tokenizer) into embedding $\mathbf{e}_{\text{spec}}$. For a batch of $B$ molecule-spectrum pairs, the model computes a similarity matrix scaled by temperature $\tau_{\text{cl}}$:
@@
where $\mathbf{E}_{\text{mol}}\in\mathds{R}^{B\times d}$ and $\mathbf{E}_{\text{spec}}\in\mathds{R}^{B\times d}$ are the normalized embedding matrices. The contrastive loss is computed bidirectionally:
@@
Molecular generation loss of NMRPeak-G. NMRPeak-G performs conditional molecule generation from spectral information. The model takes as input spectral tokens $\mathbf{z}$ and optionally molecular formula tokens $\mathbf{f}$ from the chemistry-aware adaptive tokenizer, and autoregressively generates SMILES tokens $\{s_{1},s_{2},\ldots,s_{L}\}$ obtained from Smirk-based tokenization [wadell2024smirk]:
@@
5.4.2 Model architecture
The overall framework is built upon two backbone architectures: a Uni-Mol [zhouuni] encoder for molecular representation learning and a BART-based [lewis2020bart] encoder–decoder architecture for SMILES sequence and spectral sequence modeling. Depending on the task, different combinations are employed while maintaining a unified design philosophy.
For all modules involving molecular structure encoding, we adopt the Uni-Mol encoder architecture [zhouuni]. The input consists of atomic element types and 3D coordinates, following the original Uni-Mol preprocessing protocol. The pretrained Uni-Mol weights are initialized from the official checkpoint but are not frozen; instead, the entire encoder is fine-tuned during training to adapt to NMR-related tasks. For modules involving molecular sequence generation, SMILES tokens obtained via Smirk-based tokenization [wadell2024smirk] are processed using a BART decoder. The decoder follows the standard BART decoder configuration and autoregressively generates SMILES tokens conditioned on encoder outputs. Spectral tokens produced by the chemistry-aware adaptive tokenizer are processed using a standard BART architecture. Specifically, the spectral encoder and spectral decoder both follow the BART-base configuration. All BART-based components share the same vocabulary. When encoder and decoder are both from BART, their token embedding layers are shared, consistent with the original BART design.
Architecture of NMRPeak-P. NMRPeak-P adopts a molecule-to-spectrum encoder–decoder architecture. The Uni-Mol encoder encodes molecular features into continuous representations, which are then fed into a BART decoder that autoregressively generates spectral tokens discretized by the chemistry-aware tokenizer.
Architecture of NMRPeak-R. NMRPeak-R employs a dual-encoder architecture for contrastive learning. Molecular structures are encoded by the Uni-Mol encoder, while spectral tokens are encoded by a BART encoder. To align the representation spaces, a fully connected projection layer (adapter) is appended to the Uni-Mol encoder output, projecting molecular embeddings to the same dimensionality as the spectral encoder (768 dimensions). The resulting molecular and spectral embeddings are $\ell_{2}$-normalized before computing the similarity matrix for contrastive training. In this module, no decoder is used.
Architecture of NMRPeak-G. NMRPeak-G follows a spectrum-to-molecule encoder–decoder architecture. Spectral tokens (and optionally molecular formula tokens) are first encoded by a BART encoder. The encoded spectral representations are then consumed by a BART decoder that autoregressively generates SMILES tokens using Smirk-based tokenization [wadell2024smirk]. Since both encoder and decoder follow the standard BART configuration, they share token embedding parameters. When molecular formula tokens are provided, they are concatenated with spectral tokens at the encoder input level, enabling the decoder to condition generation on both spectral patterns and elemental composition constraints.
5.4.3 Multi-model ensemble strategy for NMRPeak-P
For inference with a single-model version of NMRPeak-P, we use the best-performing checkpoint selected according to the validation performance. For the multi-model ensemble version, we leverage multiple checkpoints saved at different training steps. Instead of relying on a single best model, predictions are generated from several intermediate checkpoints and then aggregated during inference. By default, we ensemble 10 models selected from different training steps.
5.4.4 Retrieval weight calibration for NMRPeak-R
The combined score is computed as a linear combination of SME, SSE, and SSR:
$S_{\text{combine}}=\alpha\cdot\text{SME}+\beta\cdot\text{SSE}+\gamma\cdot\text{SSR},$ (26)
where $\alpha+\beta+\gamma=1$. Each coefficient takes values in the range $[0,1]$ with an interval of 0.1. Therefore, for each fixed input condition, there are 66 possible coefficient combinations.
When the input simultaneously contains both 13C and 1H spectra, SME and SSE are computed directly on the combined input, yielding a single score for each metric. In contrast, the SSR score is calculated as a weighted combination of the two modalities:
$\text{SSR}=\lambda\cdot\text{SSR}_{{}^{13}\mathrm{C}}+(1-\lambda)\cdot\text{SSR}_{{}^{1}\mathrm{H}},$ (27)
where $\lambda\in[0,1]$ with a step size of 0.1, and the two weights sum to 1.
5.4.5 Beam search strategy for NMRPeak-G
The beam search implementation follows the decoding strategies used in XLM models [conneau2019cross] and the NAG2G model [yao2024node]. During inference, we adopt task-specific decoding hyperparameters to balance determinism and diversity. For NMRPeak-P, the default beam size is set to 1 and the temperature is set to 1.0, resulting in deterministic and stable spectral predictions. For NMRPeak-G, the default beam size is set to 10 and the temperature is set to 3.0, encouraging higher diversity in generated molecular structures.
5.4.6 Computational resources
All models were trained on NVIDIA A100 GPUs with 80GB memory. NMRPeak-P and NMRPeak-G were trained on 8 $\times$ A100 80GB GPUs using distributed data parallel training, whereas NMRPeak-R was trained on a single A100 80GB GPU.
5.5 Evaluation protocol
The evaluation of NMRPeak-P is conducted from two complementary perspectives. First, we compute the similarity score using the peak-aware similarity metric, which directly measures the agreement between predicted and reference spectra at the peak level (including chemical shifts, multiplicities, integrations, and coupling constants). Notably, this metric is highly stringent, and even minor deviations in peak attributes can lead to noticeable score reductions. As a result, similarity scores are typically below 0.95, even for high-quality predictions. Second, we assess the structural consistency of predicted spectra by feeding them into a frozen NMRPeak-G model to reconstruct molecular structures. We report top-$k$ accuracy, measuring whether the ground-truth molecule appears among the top-$k$ generated candidates. This evaluation reflects the practical utility of predicted spectra in downstream structure elucidation. Due to the strictness of the peak-aware similarity metric, even small numerical differences between NMRPeak-P-Single and NMRPeak-P-Multi correspond to substantial differences in actual spectral parsing accuracy.
For NMRPeak-R and NMRPeak-G, the evaluation metrics are based on top-$k$ retrieval and generation accuracy, respectively. Notably, the retrieval accuracy of NMRPeak-R is reported over the entire test set as the candidate search space.